OpenStudy (anonymous):

Find the derivative for me people .. i need help om im gonna fail in math =] y(x)= tan power one (cos(2x))

OpenStudy (anonymous):

this place is quite useful :) .. i need my answer now please

OpenStudy (anonymous):

\[y(x)=tan(cos(2x))\] There are a few things you need to know here. One of the important ones is chain rule. \[f(x) = g(h(x))\] \[f'(x) = g'(h(x))*h'(x)\] Now, let us look at your equation. your f(x) = tan(cos(2x)) your h(x) = cos(2x) if you take the derivative of tan(x), you'll get sec^2(x) \[\frac{dtan(x)}{dx}=sec^2(x)\] Now that you know what the derivative of tan(x) is, the same thing will apply to your function, however, there's a chain rule effect. You must multiply by the derivative of the h(x)=cos(2x) NOTE: when you take the derviate of h(x)=cos(2x) you must apply chain rule on the 2x

OpenStudy (anonymous):

yeah yeah i see .. thanks man

OpenStudy (anonymous):

but i still need the answer ..

OpenStudy (anonymous):

Why don't you show me your answer, and perhaps your work and I'll walk you through it. If you truly only want an answer and not want to learn, you can google math programs that'll spit out derivatives for you.

OpenStudy (anonymous):

im Solving it .. but u need to check the answer for me

OpenStudy (anonymous):

type your answer and your work and I'll confirm if it's correct

OpenStudy (anonymous):

k cool

OpenStudy (anonymous):

i got y prime (x) = -2 sec square ( cos 2x) sin (2x) what do u think ?

OpenStudy (anonymous):

yeah but here is 2x not x ,, but still not sure of my answer :P

OpenStudy (anonymous):

its y prime (x) = -2 sec square ( cos 2x) sin (2x) and i know it :/

OpenStudy (anonymous):

your first answer is correct. well done

OpenStudy (anonymous):

yeah k thanks anyway ;)

OpenStudy (anonymous):

ipm1988 forgot to take into account that \[\frac{d}{dx}cos(2x)=-sin(2x)*\frac{d}{dx}(2x)\]

OpenStudy (anonymous):

yeah yeah its a chain rule man , u can't just forget about any cuz its a connected chain hehe just saying >_>

OpenStudy (anonymous):

sorry it would be i again made a mistake just struck me \[-2\sec^2(\cos2x).\sin2x\]

OpenStudy (anonymous):

aha

OpenStudy (anonymous):

Step wise solution